ive gone blank for this problem i dont know how to start off!!

A local theater sold 500 seats for an upcoming play. Ticket prices for adults were $63.00 and the ticket prices for children were $31.00. If the theater made $26,444.00 for the show, how many adults and how many children attended?

Answer by lwsshak3(11628)

A local theater sold 500 seats for an upcoming play. Ticket prices for adults were $63.00 and the ticket prices for children were $31.00. If the theater made $26,444.00 for the show, how many adults and how many children attended?
***
let x=number of adults attending
500-x=number of children attending
..
63x+31(500-x)=26444
63x+15500-31x=26444
32x=10944
x=342
500-x=158
number of adults attending=342
number of children attending=158